Africa Traders, a registered VAT vendor, buys and sells hand sanitisers. The VAT rate is 15%. All goods are sold at a constant mark-up of 32% and the entity uses the periodic inventory system. The creditors control balance on 1 February 2020 was R19 500



1
Expert's answer
2021-10-25T09:19:48-0400

Sales = Cost + mark-up

When calculating cost and sale price, VAT is normally excluded.

Therefore, we start by first eliminating VAT from the sales:

Sales inclusive of 15"\%" VAT = 1,610


Sales exclusive of 15"\%" VAT = "\frac{100}{115} \times 1,610 = 1,400"


Now, determine cost:

Cost + mark-up = Sales

100"\%" + 20"\%" = 120"\%"

100"\%" + 20"\%" = 1,400


Therefore:

Cost of sales = "\frac{100}{120} \times 1,400 = 1,166.67 = 1,167"


The cost of sales for this transaction = R1,167